2(3x+4)=x+6
6x+8=x+6
6x-x=6-8
5x=-2
x=-2/5

Answer by chen.aavaz(62)

2(3x+4)=x+6
6x+8=x+6
6x-x=6-8
5x=-2
x=-2/5